A plane flies 900 miles with a tail wind in 3 hours. It takes the same plane 4 hours to fly the 900 miles when flying against the wind. What is the plane's speed still in the air?
Answer by checkley71(8403)

900/3=300
900/4=225
300-225=75
75/2=37.5 SPEED OF THE WIND.
300-37.5=262.5 SPEED OF THE PLANE IN STILL AIR.
-------------------------------------------------------
If you should have time there is 1 other problem I am stuck on here it is Kathy and Brian drove a total if 674 miles in 111.4 hours. Kathy drove the first part of the trip and averaged 65 miles per hour. Brian drove the remainder of the trip and averaged 55 miles per hour. For what length of time did Kathy drive?
BELIEVE THE 111.4 HOURS IS INCORRECT.
674/111.4=6.05 MPH WOULD BE THE AVERAGE. HOWEVER THE ANSWER MUST BE BETWEEN 55 & 65 MPH.
ASSUMING THE HOURS=11.14 THEN:
D=RT OR T=D/R
X/65+(674-X)/55=11.14
[55X+65(674-X)]/65*55=11.14
[55X+43,810-65X]/3,575=11.14
55X-65X+43,810=11.14*3,575
-10X=39,825.5-43,810
-10X=-3,984.5
X=-3,984.5/-10
X=398.45 MILES IS THE DISTANCE DRIVEN @ 65 MPH.
674-398.45=275.5 FOR THE 55 MPH DRIVE.
PROOF:
398.45/65+275.5/55=11.14
6.13+5.01=11.14
11.14=11.14
Thank you again for everything!
Cheryl